eBook Elliott, Inc, has four salaried clerks to process purchase orders. Each clerk is pald a salary of $26,800 and is capable of processing as many as 6,550 purchase or printer in processing orders. Time available on each PC system is sufficient to process 6,550 orders per year. The cost of In addition to the salaries, Ellittspends $28,296 for forms, postage, and other supplies (assuming 26,200 PC system is $1,000 per year. the salaries, processed). During the year, 25,500 orders were processed Elliott spends $28,296 for forms, postage, and other supplies (assuming 26,200 purchase orders are Required: 1. Which of the resources associated with purchasing can be classified as flexible resources? Forms, Postage and Supplies Which of the resources associated with purchasing can be classified as committed resources? PC system and clerks 2. Compute the total activity availability, and break this into activity usage and unused activity. Activity availability: 26,200 purchase orders Activity usage: 25,500 purchase orders Unused activity: 700 purchase orders and the cost of unused activity. When required, roundy late the total cost of resources supplied (activity cost), and break this into the cost of activity used decimal places. answers to the nearest dollar. If required, round your intermediate calculations to five s138,740 Activity cost: Cost of activity used: Cost of unused activity:
Answer #1

Solution 1:

Forms, Postage and supplies resources associated with purchasing canbe classified as flexible resources.

Solution 2:

Activity availability = 26200 Purchase order

Activity usage = 25500 purchase order

Unused activity = 26200 - 25500 = 700 purchase orders.

Soltuion 3:

Total cost of rescources supplied = Committed cost + Flexible cost

Committed cost = Salary of clerk + Cost of PC system = $26,800*4 + $1,000*4 = $111,200

Flexible cost = Cost of forms postage and other supplies = $28,296 / 26200 * 25500 = $27,540

Total activity cost = $111,200 + $27,540 = $138,740

Cost of activity used = cost of forms postage and other supplies + Proportinate committed cost based on total activity

= $27,540 + $111,200 * 25500 / 26200 = $135,769

Cost of unused activity = $138,740 - $135,769 = $2,971